### v1

# DamageCAT: A Deep Learning Transformer Framework for Typology-Based Post-Disaster Building Damage Categorization

Abstract
Natural disasters increasingly threaten communities worldwide, creating an urgent need for rapid, reliable building damage assessment to guide emergency response and recovery efforts. Current methods typically classify damage in binary (damaged/undamaged) or ordinal severity terms, limiting their practical utility. In fact, the determination of damage typology is crucial for response and recovery efforts. To address this important gap, this paper introduces DamageCAT, a novel framework that provides typology-based categorical damage descriptions rather than simple severity ratings. Accordingly, this study presents two key contributions: (1) the BD-TypoSAT dataset containing satellite image triplets (pre-disaster, post-disaster, and damage masks) from Hurricane Ida with four damage categories (partial roof damage, total roof damage, partial structural collapse, and total structural collapse), and (2) a hierarchical U-Net-based transformer architecture that effectively processes pre-post disaster image pairs to identify and categorize building damage. Despite significant class imbalances in the training data, our model achieved robust performance with overall metrics of 0.7921 Intersection over Union (IoU) and 0.8835 F1 scores across all categories. The model’s capability to recognize intricate damage typology in less common categories is especially remarkable. The DamageCAT framework advances automated damage assessment by providing actionable, typological information that better supports disaster response decision-making and resource allocation compared to traditional severity-based approaches.
keywords:
Damage assessment, Satellite imagery, Transformers, Damage description
1 Introduction
1.1 Motivation & problem statement
The increasing frequency and severity of natural disasters pose significant challenges to emergency response systems worldwide. According to Gholami et al. [1], natural disasters affect approximately 350 million people annually, causing billions of dollars in damages. Climate change is expected to further exacerbate these trends, leading to more frequent and intense disasters [2]. In this context, rapid and accurate building damage assessment becomes crucial for effective disaster response and recovery. Building damage assessment serves multiple critical purposes in disaster response. As Adriano et al. [3] emphasizes, timely damage assessment is essential for both emergency response and prompt recovery efforts. Similarly, Dias et al. [4] highlights that building damage assessment is vital for guiding disaster response missions and estimating recovery needs across impacted areas. The information gathered through damage assessment enables humanitarian organizations to be prompt, effective, and efficient in supporting affected populations [5]. Traditional damage assessment methods rely heavily on manual visual interpretation of post-disaster imagery. However, this approach presents significant limitations. Alisjahbana et al. [6] point out that manual interpretation is not only time-consuming but often results in very low accuracy. The labor-intensive and tedious nature of manual annotation further hampers rapid response efforts [5]. In addition, field-based assessments may be dangerous or impossible in severely affected areas, creating information gaps in critical locations. Satellite imagery offers a promising alternative for building damage assessment. In recent years, the use of remote sensing data and machine learning has significantly enhanced damage assessment in disasters. Kahl and Chen [7] note that satellites provide a unique, comprehensive viewpoint for information retrieval from disaster-affected zones. Commercial satellite providers now enable near-daily monitoring of Earth’s surface, facilitating timely post-disaster assessments [4]. Moreover, remote sensing technologies provide rich and reliable information to support expert decision-making [5]. Despite these advantages, satellite-based damage assessment faces several challenges. Chen et al. [8] observed that the task is complicated by diverse building structures and complex environments. Liu et al. [9] highlight that building damage classification is challenging due to limited publicly available imagery data of damaged buildings and the difficulty in judging damage levels from images. Furthermore, most existing deep learning models are designed for non-damaged building segmentation and may not be suitable for damaged building classification [9]. The integration of deep learning with satellite imagery has shown promising results in automating building damage assessment. Neto and Dantas [10] demonstrated the potential of combining mathematical morphology and convolutional neural networks for building damage segmentation. Similarly, May et al. [11] utilized advanced networks such as EfficientNet and Siamese models to detect buildings and evaluate damage levels using pre- and post-event images. However, several limitations persist in current deep-learning approaches for remote sensing-based disaster-damaged assessment. Melamed et al. [12] identified bias in commonly used datasets, noting that deep learning models struggle to accurately identify isolated damaged buildings, potentially causing oversights in critical disaster scenarios and delaying humanitarian aid. In addition, Wang et al. [13] point out challenges related to extremely imbalanced datasets, where damaged buildings are significantly underrepresented compared to undamaged ones. Another significant limitation of current deep learning-based models for building damage assessment is their reliance on subjective severity scales (e.g., “low,” “medium,” “high”), which offer limited insight into the specific nature and type of damage. These severity labels often arise from coarse-grained assessments that can vary between datasets or annotators, leading to inconsistencies and ambiguity. Moreover, a single severity score cannot distinguish whether damage is confined to a roof or extends into load-bearing structures, nor can it differentiate between partial and total collapse. As a result, key nuances—such as the potential for secondary hazards or the requisite expertise for repairs—are frequently overlooked. This lack of descriptive granularity ultimately hampers targeted resource allocation, informed decision-making, and effective prioritization of rescue and recovery efforts in post-disaster scenarios.
Despite the growing body of literature leveraging deep learning for post-disaster damage assessments, current severity-based classification approaches often provide only coarse estimates (e.g., “low,” “moderate,” or “high” damage). Such subjective rating scales neglect the underlying mechanisms or precise nature of the damage, limiting the applicability of these models for decision-making in real disaster scenarios. Knowing whether a structure exhibits partial roof damage or total structural collapse is not just a matter of granularity; it directly influences emergency response strategies, resource allocation, and subsequent recovery efforts. For instance, responding to a partially damaged roof may call for specialized roofing repair teams and temporary protective measures, whereas a structural collapse demands immediate search-and-rescue operations and comprehensive rebuilding plans. This study addresses that critical information gap by focusing on descriptive, type-based classifications that capture the distinct modes of structural failure, thereby enhancing both the accuracy and practical utility of automated damage assessment systems. This research context emphasizes the demand for improved approaches that reliably evaluate building damage typology across various disaster conditions, construction types, and geographic regions.
1.2 Research objectives
This study aims to advance building damage assessment by moving beyond traditional severity labels to a typology-based classification framework that captures distinct modes of structural failure. Specifically, we present a deep learning architecture that processes pre- and post-disaster satellite imagery to identify and categorize building damage with greater detail. To achieve this, we introduce BD-TypoSAT (Building Damage Typology Satellite Dataset), a novel resource containing four specific damage categories—partial roof damage, total roof damage, partial structural collapse, and total structural collapse—offering richer representation than standard severity-based labels. We then deploy a transformer-driven, hierarchical U-Net that integrates semantic segmentation and change detection to precisely segment building footprints and classify them by damage type. This approach yields actionable information for emergency managers and stakeholders. Finally, we rigorously evaluate the model’s performance using multiple metrics (e.g., IoU, F1) to demonstrate its utility under real-world conditions, and we discuss the operational implications for disaster response, such as more informed resource allocation and prioritization of recovery efforts. By delivering detailed damage classifications—rather than broad severity ratings—our framework aims to enhance the accuracy, utility, and timeliness of post-disaster building assessments, ultimately strengthening emergency response and community resilience.
1.3 Contributions
This study makes several key contributions that collectively enhance the field of automated building damage assessment. First, we introduce the BD-TypoSAT dataset, a human-annotated resource derived from high-resolution satellite imagery captured after Hurricane Ida (2021), which departs from conventional repositories by defining four distinct damage types—partial roof damage, total roof damage, partial structural collapse, and total structural collapse—instead of relying on binary or coarse severity labels. Second, we present a novel hierarchical U-Net-based transformer model designed to process pre- and post-disaster images and directly output typological damage classifications, offering granular and actionable insights rather than generic high/medium/low ratings. Third, we tackle the class imbalance endemic to real-world disaster data by implementing weighted loss functions (combining cross-entropy and Dice loss) and data augmentation strategies, achieving robust classification performance with an overall IoU of 0.7921 and F1 of 0.8835. In doing so, the framework provides targeted decision support, distinguishing between roof-focused damage and deeper structural failures, which ensures more effective resource allocation than severity-based methods that can often be ambiguous. Furthermore, the empirical validation on Hurricane Ida data underscores the model’s feasibility and scalability, even in complex urban environments. Lastly, by releasing both the BD-TypoSAT dataset and our transformer-driven architecture, this research lays a foundational platform for future studies aiming to expand typological damage classification to other hazards or incorporate additional data modalities, ultimately revolutionizing how post-disaster assessments and recovery efforts are conducted.
Also, this study offers novel computational contributions by integrating a hierarchical U-Net-based transformer architecture with a multi-scale difference-block mechanism that jointly processes pre- and post-disaster satellite images for fine-grained damage detection. Unlike traditional CNN pipelines, the transformer backbone facilitates long-range contextual learning, while the U-Net structure preserves critical spatial detail across multiple resolution levels. At each stage, difference blocks explicitly compare corresponding features from the two temporal images, enabling the model to pinpoint even subtle structural changes. To combat the severe class imbalance inherent in real disaster data, a weighted combination of cross-entropy and Dice losses ensures that rarer damage categories—such as total structural collapse—are not overlooked. This synergy of hierarchical feature extraction, transformer-based global context, and tailored loss weighting delivers robust, typology-specific classification performance, pushing beyond binary or severity-only assessments toward a more detailed, actionable framework for post-disaster damage analysis.
1.4 Paper organization
First in Section 2, we review some previous work that is highly related to this paper. Later we introduce the method, including the dataset, experimental setup, and the model training details, in Section 3. In Section 4, we discuss the proposed framework’s significance and list limitations and challenges. Finally, we conclude the work and give future research suggestions in Section 5.
2 Related work
Research on automated building damage assessment has rapidly evolved, leveraging a wide variety of deep learning, remote sensing, and GIS-based techniques. Researchers use change detection and temporal fusion to capture pre- and post-disaster variations. For example, Zheng et al. [14] presented a deep object-based semantic change detection framework that accelerates disaster response by directly comparing pre- and post-event images. Ismail and Awad [15] proposed BLDNet, a semi-supervised framework that uses graph convolutional networks and urban domain knowledge to reduce labeling costs, even if its reliance on graph-based relationships may falter in areas with atypical urban layouts. Weber and Kané [16] fused multi-temporal satellite imagery to enhance damage assessment in dynamic environments, even though inconsistent image quality sometimes degrades its accuracy. Zhan et al. [17] employed a Siamese convolutional network to match pairwise images, boosting discriminative power while suffering when spatial alignment falters. Chen et al. [18] introduced ChangeMamba, which formulates change detection via a spatiotemporal state space model that manages uncertainties in temporal sequences, yet tuning its parameters on large datasets remains challenging. Ahn et al. [19] combined a vision foundation model with change detection to improve generalizability across disasters, although it struggles with highly specialized local scenarios. Mohammadian and Ghaderi [20] proposed Siamixformer, a fully transformer-based Siamese network that fuses temporal features for fine-grained extraction, while its high memory consumption limits its real-time application.
Many studies enhance network architectures to better address damage assessment tasks. Gu et al. [21] surveyed rapid building damage identification methods based on remote sensing, emphasizing the efficiency of CNNs and transformer-based methods while noting that a lack of standardized datasets leads to inconsistent performance comparisons. Oludare et al. [22] proposed an attention-based two-stream high-resolution network tailored for building damage assessment from satellite images. Their method fuses high-resolution spatial features with attention modules to achieve strong damage identification accuracy, yet its heavy reliance on large labeled datasets limits adaptability in regions with sparse annotations. Shen et al. [23] introduced BDANet—a multi-scale CNN with cross-directional attention that effectively captures both local and global damage cues, though its high computational overhead makes it less suited for resource-constrained settings. Roohi et al. [24] combined multiple classifiers in an ensemble machine learning approach to improve flood damage predictions, even though the method requires careful tuning of numerous hyperparameters. Chen et al. [25] built a dual-task Siamese transformer framework that tackles segmentation and damage classification simultaneously; this joint optimization yields richer features but increases computational demands. Kaur et al. [26] designed a hierarchical transformer that organizes features at local, regional, and global scales to improve large-scale assessment, yet balancing depth and inference speed remains difficult. Gupta and Shah [27] introduced RescueNet, which unifies building segmentation and damage assessment into a single network to streamline training; however, the model sometimes confuses segmentation with damage levels. Some research [28, 29, 30], contributed foundational architectural modules—such as Squeeze-and-Excitation blocks, U-Net, and UNet++—that many damage assessment methods incorporate. Although these innovations boost feature recalibration and segmentation performance, they add extra parameters and memory requirements that complicate deployment in resource-limited settings.
Researchers also innovate in training approaches and dataset creation to overcome the shortage of labeled data. Miglani and Kumar [31] reviewed deep learning models for hazard-damaged building detection and synthesize various strategies, noting the difficulty in transferring models across disaster types. Singh and Hoskere [32] developed a semi-supervised transformer method that leverages ultra-high-resolution aerial imagery and unlabeled data to enhance damage detection, yet it faces heavy computational loads when processing large-scale images. Ro and Gong [33] proposed a scalable method to create annotated disaster image databases, which supports subsequent model development but may underrepresent rare disaster scenarios. Xia et al. [34] pioneered a self-supervised contrastive learning approach that reduces reliance on ground-truth labels and achieves performance on par with supervised methods, even though it struggles with fine-grained damage distinctions. Finally, Gupta et al. [35] released the xBD dataset—one of the largest repositories with over 850,000 building annotations—to serve as a benchmark for damage assessment models, despite its inherent class imbalance and image quality variability.
Some studies integrate diverse data sources to provide a more comprehensive picture of disaster impacts. Braik and Koliou [36] proposed a framework for automated damage assessment that integrates satellite imagery, GIS layers, and deep learning, highlighting the possibility of large-scale mapping. Their key contribution lies in combining multi-source data to improve building footprint extraction and damage classification, although the approach faces challenges in real-time deployment due to the complexity of multi-source data integration. Dong et al. [37] analyzed integrated infrastructure plans to enhance post-hazard access to critical facilities, bridging urban planning data with damage models, though it depends on extensive data availability. Techapinyawat et al. [38] applied deep learning with post-classification correction for urban land cover analysis, which can refine damage mapping but sometimes misses atypical disaster-induced changes. Gholami et al. [1] focus on the practical deployment of building damage assessment tools using satellite imagery, actively addressing issues like limited computing resources and data sharing constraints. Xue et al. [39] introduced a multi-modal approach that fuses street-view imagery with structured building data to produce refined damage predictions, even though its effectiveness is confined to regions with both reliable ground-level visuals and comprehensive metadata.
Despite substantial progress in satellite-based building damage assessment, most existing methods still rely on broad severity scales—such as binary “damaged/undamaged” labels or imprecise “low/medium/high” ratings—that fail to capture the complexity of real-world damage scenarios. By focusing exclusively on severity, these approaches overlook critical descriptive details, such as whether the damage is localized to the roof or extends into load-bearing structures—information that can be pivotal for determining repair priority, resource allocation, and insurance decisions. Moreover, nuanced indicators like partial collapses or roof breaches can signal specific failure mechanisms and potential secondary hazards, yet these remain unaddressed in standard severity-centric frameworks. As a result, decision-makers—ranging from emergency responders to policy planners—lack the granular insights needed to tailor interventions and expedite recovery. Addressing this shortfall necessitates integrated frameworks that move beyond measuring “how much” damage occurred to clarifying “what kind” of damage has taken place, enabling more targeted, context-aware strategies for disaster response, reconstruction, and long-term resilience planning.
3 Method
Figure 1: The overall workflow of the present study.
In this study, we build upon the hierarchical transformer architecture proposed by Kaur et al. [26] for image change detection, tailoring it to the specific requirements of post-disaster damage assessment. The model simultaneously processes pre- and post-disaster satellite images to produce a change map that highlights structural alterations between the two temporal states. Central to this approach is the difference block, a novel component designed to extract and compare multi-scale features across each stage of the network, thereby capturing subtle yet critical variations indicative of diverse damage types. This architecture effectively combines the long-range contextual learning strengths of transformers with the detailed localization capabilities of a U-Net-like design, making it particularly suited for high-resolution satellite imagery. The overall workflow is shown in Figure 1.
3.1 Dataset
Figure 2: The study area where the satellite imagery is captured.
We present the novel dataset employed for training and validating the model in this research—the BD-TypoSAT (Building Damage Typology Satellite Dataset) dataset—which was developed by us utilizing Maxar’s satellite images captured during Hurricane Ida. On Sunday, August 29, 2021, Hurricane Ida struck parts of Louisiana and Mississippi with wind gusts reaching up to 172 mph, leaving more than a million customers without electricity, including the entire New Orleans area [40]. During the disaster, Maxar captured high spatial resolution satellite imagery (at 0.4 m/pixel) and it was subsequently made publicly available [41]. Figure 2 illustrates the region where the satellite images were obtained. The original images were segmented into 512 $\times$ 512-pixel patches to maintain spatial context while enabling detailed analysis. From this process, we generated a dataset of 2,135 triplets, each containing a pre-disaster image, a post-disaster image, and a manually annotated damage categorical mask.
Figure 3: Example images with manually labeled annotations; the legend of the four damage types are shown below the picture.
In this study, building damage is categorized into four specific classes—partial roof damage, total roof damage, partial structural collapse, and total structural collapse—thereby capturing key structural distinctions that go beyond coarse severity labels. Unlike many comparable datasets, our annotation exclusively focuses on damaged structures, deliberately omitting undamaged buildings to ensure that all labeled instances represent actual disaster-induced harm. As illustrated in Figure 3, each damage type is color-coded for clarity, highlighting the diverse manifestations of structural failure. However, the distribution of annotated buildings skews toward less severe damage, making the rare instances of severe collapse comparatively underrepresented. To address this class imbalance during training, we adopt a composite loss function that fuses weighted cross-entropy with Dice loss, assigning each category a weight derived as the square root of its inverse frequency (see Table 1). By boosting the influence of scarce yet critical classes, this weighting strategy ensures that the model remains sensitive to all damage types, including those often overshadowed by more prevalent, lower-severity categories.
Damage description Number of buildings Percentage Weight in loss function (Square root of inverse frequency)
Partial roof damage 7030 86.69% 1.07
Total roof damage 958 11.81% 2.91
Partial structural collapse 83 1.02% 9.9
Total structural collapse 38 0.47% 14.61
Table 1: Distribution of building damage categories with corresponding frequency weights in loss function
The model employs a composite loss function formulated as:
$Loss=\sum_{n\in C}w_{n}(0.5\cdot dice_{n}+0.5\cdot ce_{n})$ (1)
where $C$ represents the set of all categories, $w_{n}$ denotes the weight assigned to the $n$-th category, $dice_{n}$ is the dice loss, and $ce_{n}$ is the cross entropy loss for category $n$. Leveraging both Dice loss and cross-entropy in a composite loss function addresses the respective limitations of using either metric in isolation. Dice loss places greater emphasis on the spatial overlap between predicted and ground-truth masks, making it particularly effective in highly imbalanced segmentation tasks where underrepresented classes might otherwise be neglected. However, Dice loss can suffer from unstable gradients when the overlap is minimal—especially in early training or in scenarios with rare damage categories. In contrast, cross-entropy excels at pixel-wise classification accuracy, but it can struggle to account for severe class imbalances, potentially skewing model performance toward the majority class. By combining these two losses, we harness Dice loss’s ability to robustly guide the segmentation of underrepresented classes while benefiting from cross-entropy’s stable gradient flow and accurate pixel-level learning. This unified approach thus ensures balanced recognition of each damage category, from minor roof impairments to total structural collapse.
3.2 Model Architecture
Figure 4: Transformer and U-Net-based architecture for disaster change detection and segmentation
Figure 4 presents the DAHiTrA model we are using in this work, which is proposed by Kaur et al. [26]. This hierarchical U-Net architecture is designed for satellite imagery change detection in disaster scenarios. The network processes pre-disaster and post-disaster images ($3\times 512\times 512$) through parallel ResNet-based down-sampling paths that progressively extract hierarchical features at decreasing resolutions. Each green block represents a down-sampling stage that halves spatial dimensions while increasing feature depth, creating multi-scale feature representations from $64\times 256\times 256$ down to $256\times 32\times 32$. At three intermediate levels, the difference blocks compute and analyze changes between corresponding feature maps from both temporal states. These difference blocks capture change information at multiple scales, from fine details to broader contextual changes. The decoders (purple blocks) include a dual convolutional block with intermediate normalization and 2×2$\times$2 × up-sampling operation to gradually restore spatial resolution while integrating change information from multiple scales. Skip connections (dashed lines) provide direct pathways for high-resolution spatial information to flow into the decoder. The network ends with a concatenation operation that merges the highest-level features with the original resolution features. This creates a final $5\times 512\times 512$ output that classifies each pixel into five distinct change categories (one of which is the background). This architecture effectively balances fine-grained spatial detail with hierarchical feature representation, making it suitable for identifying structural changes in post-disaster environments.
3.3 Experimental Setup
All experiments were conducted on a workstation equipped with an NVIDIA RTX 6000 GPU with 48GB of VRAM. We trained our model using the following hyperparameters: learning rate of 0.001 with a linear decay policy, input image size of 512 $\times$ 512 pixels, batch size of 8, and a maximum of 300 epochs. The Adam optimizer was used with default $\beta$ values ($\beta_{1}=0.9$, $\beta_{2}=0.999$) and $\epsilon=10^{-8}$. Early stopping was implemented with a patience parameter of 20 epochs to prevent overfitting, and the model weights from the epoch with the lowest validation loss were selected for evaluation. We implemented a custom data augmentation pipeline to increase the effective size of the dataset. Images were processed at 512 $\times$ 512 pixels and normalized with mean and standard deviation values of [0.5, 0.5, 0.5]. Our augmentation strategy included random horizontal and vertical flipping, 90/180/270 rotations, and Gaussian blur with random radius. During training, we employed a patch-based cropping approach where random crops were extracted from training images, while validation used predefined patch positions. To address the class imbalance issue, we implemented an upsampling technique that duplicated training samples containing specific class labels (particularly those with positive values in classes 2-4), with additional duplication for samples containing classes 2 or 3, effectively increasing the representation of less frequent classes in the training dataset. The model was implemented using PyTorch 2.5.1.
3.4 Quantitative Evaluation
Figure 5: Validation metrics during the training procedure.
We evaluated the model’s performance using a variety of metrics, including Intersection over Union (IoU), F1 score, precision, and recall. Table 2 summarizes the performance metrics for each damage category, as well as the overall performance across all categories. The model was trained for the full 300 epochs as specified in our experimental setup. We monitored several key metrics throughout the training process, including training loss, validation loss, IoU (Intersection over Union), and F1-score. Figure 5 illustrates the progression of various performance metrics throughout the training process over 300 epochs. The overall metrics (Figure 5 (a)) demonstrate a consistent upward trajectory, with Mean F1 score showing more rapid initial improvement compared to Mean IoU, ultimately converging to approximately 0.883 and 0.792, respectively. The training and validation loss curves (Figure 5 (c)) show typical decreasing patterns, stabilizing after 150 epochs. The minimal gap between curves indicates good generalization without overfitting.
Class-specific performance metrics reveal notable insights into the model’s ability to distinguish among the four damage categories. In particular, the partial structural collapse class attained the highest overall performance (IoU $\approx 0.839$, F1 $\approx 0.876$ at epoch 298), albeit with significant fluctuations between epochs 25 and 100. This strong yet variable outcome may stem from the visually distinctive cues associated with partial structural failure—such as exposed interior walls or partial debris fields—making it easier for the model to recognize, even if the early training stages faced instability. Meanwhile, both roof damage categories (partial and total) produced comparable IoU scores (ranging from 0.74 to 0.82) and F1 scores (0.85 to 0.90), despite having the largest number of annotated instances. This paradox suggests that visual similarity—for example, two roofs both missing shingles—can pose a classification challenge, underscoring how top-down satellite views often compress details. Moreover, the total structural collapse class underperformed slightly compared to its partial counterpart, largely due to data scarcity (only 38 labeled examples out of 8,109). Although our weighted loss and data augmentation strategies helped mitigate some imbalance issues, these results highlight the need for additional training samples to better capture the rare yet catastrophic nature of total collapses. Another key observation is that class-specific disparities were most pronounced before epoch 75, pointing to the longer learning curve for certain damage patterns, which require extensive iterative updates to achieve stable recognition. Nonetheless, all categories exceeded typical expectations for imbalanced datasets, achieving strong F1 scores (0.8519–0.9097). This robust performance across a range of damage types reinforces the efficacy of our modeling choices, including the difference blocks, transformer-based architecture, and balanced loss function, for capturing both the subtle and severe manifestations of disaster-induced building damage.
The class-specific disparities were most pronounced before epoch 75, suggesting that certain damage patterns required more training iterations to be accurately recognized. Despite implementing various techniques to address the class imbalance issue, acquiring additional samples from the underrepresented class remains necessary. Overall, contrary to typical expectations in imbalanced datasets, all categories demonstrate strong performance with F1 scores ranging from 0.8519 to 0.9097.
Category IoU F1 Precision Recall
Partial roof damage 0.7420 0.8519 0.8759 0.8292
Total roof damage 0.8128 0.8967 0.9357 0.8609
Partial structural collapse 0.8344 0.9097 0.9332 0.8874
Total structural collapse 0.7790 0.8758 0.9489 0.8132
Overall 0.7921 0.8835 0.9234 0.8477
Table 2: Performance metrics for each damage category
3.5 Qualitative Analysis
To supplement the quantitative metrics, we conducted a visual evaluation of model performance using five representative image samples, selected to encompass diverse damage scenarios and varied building typologies (Figure 6). Each sample contains pre- and post-disaster satellite views alongside the model’s prediction mask and ground truth annotations, offering a side-by-side comparison of detection and classification outcomes.
This qualitative examination reveals a high degree of concordance between predicted masks and annotated references, underlining the model’s accuracy in distinguishing multiple damage categories. Even in dense urban districts—where numerous damaged structures can sit in close proximity—the model effectively delineates individual building boundaries and assigns accurate damage labels, demonstrating fine-grained spatial discrimination. Notably, it consistently recognizes differences between partial roof damage and partial structural collapse in adjacent buildings, reflecting a nuanced grasp of visual cues associated with each damage type.
Overall, these visual findings underscore the model’s practical utility in emergency response contexts, where rapid, detailed mapping of post-disaster impacts is essential. Unlike binary or severity-only classification methods, this approach delivers categorical insights—pinpointing whether a roof is simply compromised or an entire structure is at risk of collapse. Such richer information can prove invaluable in guiding targeted interventions and resource allocation. Moreover, the consistency of these results across varied scenarios validates the robustness of our typology-based system, especially given the novelty of defining damages with such granular specificity and the absence of direct benchmarks in the current literature.
Figure 6: The qualitative evaluation result of 5 select image data points.
4 Discussion
4.1 Interpretation of results
Our experimental findings underscore the viability of a categorical, deep-learning-driven approach to building damage assessment, leveraging pre- and post-disaster satellite imagery. The performance metrics for the four damage categories provide important insights into the method’s strengths and remaining challenges. Notably, the rare “partial structural collapse” class achieved unexpectedly strong results (IoU $\approx 0.8344$, F1 $\approx 0.9097$), surpassing metrics for the more common “partial roof damage” category (IoU $\approx 0.7420$, F1 $\approx 0.8519$). This discrepancy highlights that visual distinctiveness (e.g., exposed interior walls or collapsed debris fields) can override class imbalance constraints, making certain damage patterns more readily identifiable from a top-down viewpoint.
However, persistent performance gaps across categories—even after applying weighted losses and data augmentation—emphasize the intrinsic difficulties in using vertical satellite imagery for detailed damage classification. Roof damage, in particular, presents subtle visual cues (e.g., missing shingles or partial roof tarps) that can be masked by factors like shadowing, overhanging vegetation, or variations in roof color. Consequently, the two roof damage classes exhibit relatively lower IoU and F1 scores compared to structural collapse classes, whose larger-scale changes (e.g., caved-in walls, rubble) are more conspicuous in vertical imagery. Even though the 0.4,m/pixel resolution images provided valuable detail, fine-grained features remain challenging to discern due to atmospheric interference, illumination conditions, and satellite angle variations—an issue corroborated by Chen et al. [8] and Liu et al. [9].
Overall, these outcomes underscore the need for continued innovation, such as multi-modal data integration or oblique and street-level perspectives, to capture more subtle roof-related damage patterns. Nonetheless, the robust performance observed across all categories confirms the promise of our typology-based approach, offering a nuanced alternative to traditional severity-focused models and paving the way for more targeted disaster response and recovery strategies.
4.2 Significance and implications
The DamageCAT framework constitutes an important advancement by elevating building damage assessment from coarse severity ratings (e.g., “minor,” “moderate,” “major”) to typology-based classifications. This novel paradigm bridges an important research gap identified by Melamed et al. [12], enabling the deep learning community to transition from generic damage labels to specific, well-defined categories (e.g., partial roof damage vs. total structural collapse). From a methodological standpoint, our transformer-enhanced U-Net architecture effectively leverages pre- and post-disaster satellite imagery to isolate disaster-induced changes, filtering out pre-existing conditions as advocated by Zheng et al. [14] and Ahn et al. [19]. This change-focused design, coupled with multi-scale difference blocks, advances the state-of-the-art by robustly detecting subtle structural variations in highly imbalanced data settings.
Equally important are the practical implications of this typology-based approach for emergency response and recovery planning. By pinpointing whether a building has partial roof damage versus a complete structural collapse, DamageCAT furnishes actionable intelligence that drastically improves resource allocation—for instance, dispatching specialized roofing teams for moderate impairments while prioritizing heavy machinery and rescue personnel for total collapses. This level of granularity facilitates targeted interventions, reduces operational uncertainties, and potentially accelerates recovery timelines compared to generic severity ratings that merely denote “light,” “medium,” or “heavy” damage. Furthermore, the ability to differentiate among multiple damage modes enhances risk assessments, enabling insurers, city planners, and engineers to prioritize critical repairs and mitigate secondary hazards (e.g., water intrusion or imminent collapse). Such data-driven insights also lend greater transparency and consistency in post-disaster decision-making, improving situational awareness for diverse stakeholders ranging from first responders to policy makers.
4.3 Challenges and limitations
Despite the promising results, this work faces several key limitations that constrain its immediate scalability and generalizability. First, there is limited access to consistent, high-resolution pre- and post-disaster imagery pairs; obtaining uniform spatial resolution and temporal alignment can be challenging, particularly in large-scale or multi-hazard contexts. Additionally, manual annotation remains time-intensive, especially for subtle or nuanced damage categories (e.g., partial roof damage vs. partial structural collapse), which can slow dataset expansion and increase labeling costs.
From a methodological perspective, relying solely on vertical (nadir) satellite imagery can obscure certain facade-level or internal damages, leading to potential misclassifications or omissions. Moreover, our dataset exhibits significant class imbalance, most notably in the total structural collapse category (representing only 0.47% of annotated buildings), which can affect model robustness despite mitigation strategies like weighted losses. The focus on a single disaster event (Hurricane Ida) and one geographic region (Louisiana) further narrows the framework’s initial scope. As Adriano et al. [3] cautions, damage assessment methods tailored to specific locales may have limited transferability to different hazards or global settings.
Finally, computational requirements—arising from the transformer-based architecture—may hinder real-time deployment in resource-constrained environments, such as remote field operations or developing regions. In line with Gholami et al. [1], inference speed can be more critical than absolute accuracy during crises, underscoring the need for further model optimization (e.g., model distillation, quantization) to meet operational demands. Addressing these limitations through multi-modal data integration, lighter model architectures, and extensive cross-regional validation represents an important avenue for future research.
4.4 Future research directions
Future research should focus on developing balanced datasets across damage categories, disaster types, and geographic regions to enhance model generalizability. Integrating multiple data sources (e.g., oblique aerial photography, street-view imagery) to provide complementary perspectives, as suggested by Braik and Koliou [36] and Xue et al. [39], and incorporating multiple post-disaster time points to track recovery progression are also important. Furthermore, developing interpretable models that highlight visual features contributing to classifications and optimizing models for edge devices with limited computational resources are key areas for advancement. These initiatives require interdisciplinary collaboration to develop systems that provide actionable intelligence for disaster response.
5 Conclusion
This study introduced DamageCAT, a novel framework for typology-based building damage assessment that transcends the limitations of purely severity-focused methods. By proposing the BD-TypoSAT dataset—a human-annotated resource containing four distinct damage categories (partial roof damage, total roof damage, partial structural collapse, and total structural collapse)—we address the pressing need for descriptive, real-world training data that captures the multifaceted nature of post-disaster structural impairments. Concurrently, we developed a hierarchical U-Net-based transformer architecture capable of analyzing pre- and post-disaster satellite imagery to produce accurate and highly detailed damage classification maps. Through comprehensive experiments, our model achieved robust performance metrics (IoU: 0.7921, F1: 0.8835), demonstrating its ability to detect both common and rare damage types, thereby offering a powerful new tool for emergency responders.
Beyond the quantitative success, the novelty of DamageCAT lies in its paradigm shift from broad severity categories to typology-based assessment, which provides fine-grained insights essential for targeted resource allocation and actionable decision-making. By pinpointing whether a building sustains roof damage versus structural collapse, the framework fosters improved prioritization of personnel, materials, and financial support. The model’s reliance on pre-post disaster change detection also reduces false positives by filtering out pre-existing building defects, further enhancing its practical utility in time-critical scenarios.
Despite these promising results, the study faces a few key limitations. First, the dataset’s imbalance (particularly for total structural collapse) reflects the real-world rarity of certain extreme damage types, yet indicates a need for expanded, balanced datasets to fully optimize performance. Similarly, the approach currently relies on vertical satellite imagery, which may obscure critical facade-level or internal damage features. Finally, the focus on a single disaster event (Hurricane Ida) and region (New Orleans, Louisiana) limits initial generalizability, and the high computational demands of the transformer-based design may pose challenges for real-time deployment in resource-constrained contexts.
Looking ahead, future research should collect and annotate more balanced data across multiple hazards and geographies, potentially integrating multi-modal sources (e.g., oblique aerial imagery, street-view photographs, LiDAR) to broaden the model’s detection capabilities. Addressing real-time inference constraints will be critical for on-site deployment, particularly by exploring model optimization techniques (e.g., pruning, and quantization). Additionally, the development of interpretable AI methods could boost user trust and offer transparency into the visual cues the model associates with specific damage types. Such innovations, combined with cross-regional validations and in-field performance tests, will continue to refine the typology-based assessment paradigm and ensure the widespread adoption of advanced building damage classification in post-disaster management.
Future research should prioritize real-time inference for emergency deployment, integration of pre-disaster building characteristics, and field validations in diverse contexts. Additionally, using interpretable AI techniques could enhance performance and user trust by revealing the visual features influencing classification.
Declarations
[…]
Acknowledgments The authors acknowledge the work of Annie Bone, Kaleb Valentin, and Kevin Rodriguez in annotating the dataset used in this study.
[…]
Funding This material is supported by the National Science Foundation under the CAREER grant (No. 1846069). Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation.
[…]
Data and Code availability The data and code that support the findings of this study are available from the corresponding authors upon request.